OpenStudy (anonymous):

Check my answer..? Write an equation of the line passing through the point (-4,2) and is parallel to y= -x+6 I got: y= -x+6; is it correct?

OpenStudy (p0sitr0n):

Parallel line slopes are equal

OpenStudy (anonymous):

yeah i know

hartnn (hartnn):

u got same equation as given ?

OpenStudy (anonymous):

whoa.. i just realized that.. umm here hold on..

OpenStudy (anonymous):

no you would put y = -x +7

OpenStudy (p0sitr0n):

Maybe y=-x-2 ?

OpenStudy (anonymous):

i had to use this equation: y-y1=m(x-x1)

hartnn (hartnn):

start with y=-x+c put the point (-4,2) there, and find c

hartnn (hartnn):

oh, so u have m, right ? which is ?

OpenStudy (anonymous):

m=-1

OpenStudy (anonymous):

i think..?

hartnn (hartnn):

yes, because lines are parallel

hartnn (hartnn):

and (x1,y1) = ?

OpenStudy (anonymous):

(-4,2)

OpenStudy (anonymous):

so it would be: y-2=-1(x-(-4)) right?

hartnn (hartnn):

so now just put the values in y-y1=m(x-x1)

hartnn (hartnn):

yes, that is correct

hartnn (hartnn):

now simplify

OpenStudy (anonymous):

ohk so then i would distribute? y-2=-1x+4 ?

hartnn (hartnn):

u missed the brackets...

OpenStudy (anonymous):

no cuz i multiplied (x-(-4)) by -1

hartnn (hartnn):

so u get -1(x+4) = -x-4

hartnn (hartnn):

-(-4) = +4

OpenStudy (anonymous):

ohh ohk so the answer would be y=-x+-2 or y=-x-2 ?

OpenStudy (anonymous):

oh nvm

hartnn (hartnn):

yeah thats correct

hartnn (hartnn):

y=-x-2

OpenStudy (anonymous):

just keep the slope the same change the y intercepter

OpenStudy (anonymous):

ohk thank you(:

hartnn (hartnn):

welcome ^_^